sammixboo (sammixboo):

When using the equation tool, it should be using parenthesis (....) instead of these brackets [....] so it does create a new line. Just something small, but it gets annoying when creating an equation and I have to go back and change all the brackets to parenthesis, so I can have equations and words on the same line

pooja195 (pooja195):

I would agree with this sammi :)

sammixboo (sammixboo):

See I would explain a problem like this using the equation tool The first step is to subtract 2 on both sides of the equation\[\frac{ x }{ 4 } + 2 = 2\] \[\frac{ x }{ 4 } + 2 - 2= 2 - 2\] \[\frac{ x }{ 4 } = 0\] Now you multiply 4 on both sides of the equation \[\frac{ x }{ 4 } \times4 = 0\times4\] \[x = 0\]

sammixboo (sammixboo):

See how it is all spread out? When I use the parenthesis it looks like this, The first step is to subtract 2 on both sides of the equation \(\frac{ x }{ 4 } + 2 = 2\) \(\frac{ x }{ 4 } + 2 - 2= 2 - 2\) \(\frac{ x }{ 4 } = 0\) Now you multiply 4 on both sides of the equation \(\frac{ x }{ 4 } \times4 = 0\times4\) \(x = 0\)

sammixboo (sammixboo):

See how much closer and better that looks?

sammixboo (sammixboo):

It just gets on my nerve when I have to go back and edit it all :(

OpenStudy (e.mccormick):

`\[ \]` is block mode, which is standard for equations. `\( \)` is inline mode, which is standard for typed text. The equation editor was built around the idea of making equations, not inline text. Oh, and if you want to see some other differences I go over them in my tutorials, like why \(\frac{1}{x}\) and \(\dfrac{1}{x}\) are different sizes.

sammixboo (sammixboo):

I know about the dfrac :p Ya taught be about that Anyways, again, @goformit100 why are you posting that on a lot of posts?

TheSmartOne (thesmartone):

^ ikr And also I like the idea Sammi :D Have you posted it in the suggestion box? :P

OpenStudy (confluxepic):

I was thinking of the same exact idea two days ago when I was helping in math. It's really annoying with the brackets.

OpenStudy (anonymous):

@sammixboo i would believe ur doing it right :D

sammixboo (sammixboo):

No, I don't post anything in the suggestion box. I can't even open it up, sooooo

sammixboo (sammixboo):

It just loads and loads

TheSmartOne (thesmartone):

lol